# Shadow-FT: Tuning Instruct via Base

### v2

@@ B.4 Detailed table for Medical Benchmarks
AIME24 pass@1 sampling 0-shot
GSM8K(0-shot) Accuracy sampling 0-shot
GSM8K(8-shot) Accuracy sampling 8-shot
MATH Accuracy sampling 0-shot
MATH-500 Accuracy sampling 0-shot
Minerva Math Accuracy sampling 4-shot
SVAMP Accuracy sampling 0-shot
Code-3
HumanEval pass@1 sampling 0-shot
HumanEval+ pass@1 sampling 0-shot
LiveCodeBench average
- generation pass@1 sampling 0-shot
- test pass@1 sampling 0-shot
- prediction pass@1 sampling 0-shot
Reasoning-9
ARC-Challenge Accuracy sampling 0-shot
BBH(0-shot) Accuracy sampling 0-shot
BBH(3-shot) Accuracy sampling 3-shot
Drop Accuracy sampling 0-shot
GPQA Diamond Accuracy sampling 0-shot
MMLU Accuracy sampling 0-shot
MMLU Pro Accuracy sampling 0-shot
Winogrande Accuracy sampling 0-shot
TheoremQA Accuracy sampling 0-shot
Table 8: Detailed results on the math benchmarks for Table 1. Three times independent training and three times evaluation average are reported.
Model Method AIME24 GSM8K GSM8K MATH MATH Minerva SVAMP Math-7
(8-shot) (0-shot) 500 Math
@@
Method MMLU MMLU WinoG DROP ARC BBH BBH GPQA TheoremQA Reasoning-9
@@
B.5 Ranks in LoRA on Llama-3.1-8B
Figure 6: Performance tuning Llama-3.1-8B with different ranks.

### v3

@@ D.6 Weight Delta Scaling
Method MMLU MMLU WinoG DROP ARC BBH BBH GPQA TheoremQA Knowledge-9
@@
Appendix E Case Study on Llama-3.1-8B
Table 12: Case study example from MATH-500. Shadow-FT can avoid errors with the knowledge learned from the base model, while the vanilla tune method fails. We extract the key step for better visualization.
Input Question:
How many positive two-digit integers are factors of both 100 and 150?
[…]
Vanilla FT 1. **Find the GCD of 100 and 150:**
2. **Identify the two-digit factors of 50:**
stop counting at 25.
Thus, the positive two-digit integers that are factors of both 100 and 150 are 10 and 25.
[…]
Shadow-FT (Ours) ###### Step 1: Find the Greatest Common Divisor (GCD) of 100 and 150
###### Step 2: Find All Factors of the GCD (50)
###### Step 3: Identify the Two-Digit Factors
(Note: 50 is a two-digit number too, but we’ll check it as well.)
So the two-digit factors of 50 are:
\$\$10, 25, 50\$\$
###### Step 4: Count the Two-Digit Factors
###### Final Answer:
[…]
Appendix F Limitations
In Shadow-FT, we first tune the Base model and then graft the weight updates to the Instruct model. However, there are some LLMs for which the paired Base models are not available, such as Qwen3-32B and Qwen3-Next. For these LLMs, we can not apply Shadow-FT. Therefore, finding a proper ’shadow’ for these models is an interesting topic for future work.